### v1

@@ 1 Introduction
In machine learning, the idea of data manifold is based on the assumption that the data space, containing the data points on which we perform classification tasks, has a natural Riemannian manifold structure. It is a quite old concept (see [1], [2] and refs therein) and it is linked to the key question of dimensionality reduction [3], which is the key to efficient data processing for classification tasks, and more. Several geometrical tools, such as geodesics, connections, Ricci curvature, become readily available for machine learning problems under the manifold hypothesis, and are especially effective when employed in examining adversarial attacks [4, 5, 6, 7] or knowledge transfer questions, see [8, 9] and refs. therein.
More specifically, the rapidly evolving discipline of information geometry [10, 11, 12, 13] is now offering methods for discussing the above questions, once we cast them appropriately by relating the statistical manifold, i.e. the manifold of probability measures, studied in information geometry with data manifolds (see [10, 14]).
We are interested in a naturally emerging foliation structure on the data space coming through the data information matrix (DIM), which is the analog of the Fisher information matrix and in concrete experiments can be obtained by looking at a partially trained deep learning neural network for classification tasks. As it turns out, the leaves of such foliation are related with the dataset the network was trained with [14, 15]. Further work in [15] linked such study to the possible applications to knowledge transfer.
@@ 2 Preliminaries
[…]
where $s$ is called a score function, and $\mathrm{softmax}$ is the function $\left(a_{i}\right)_{i}\mapsto\left(e^{a_{i}}/\sum_{k}e^{a_{k}}\right)_{i}$. We detail why these assumptions are important in the following sections. An example of neural network satisfying these conditions is the feed-forward multi-layer perceptron with ReLU activation functions. Other popular neural network architectures verify theses assumptions.
@@
Hence, $D$ is a $d\times d$ matrix, $d$ the dimension of the dataspace.
@@ 3 Cartan moving frames
In this section we first review some basic notions of Cartan’s approach to differentiable manifolds via the moving frames point of view and then we see some concrete application to our setting.
@@ 4 The Curvature of the Data Leaves
Computing in practice the curvature of a manifold with many dimensions is often intractable though essential for many tasks. This is why we are interested in computing the curvature just for the data leaves. Notice that, since in our frame $\left(e_{k}:=\sum_{i}\partial_{i}p_{k}\left(x\right)\partial_{i}\right)$, we have $\sum_{i}\partial_{i}\ln p_{k}(x)\partial_{i}=\frac{1}{p_{k}(x)}\sum_{i}{\partial_{i}p_{k}(x)\partial_{i}}=\frac{1}{p_{k}(x)}e_{k}$, we may forget the logarithm in the computations as it contributes only by a scalar factor.
@@
The proof is a straightforward combination of Equation 9 and Equation 10. We simply need to replace $\partial_{i}p_{a}$ twice in the expression 10 by 9. ∎
@@
Then, we use the following proposition to remove the second order derivative in $M$. An alternative formula for this expression is given in A.1. This facilitates the computations with automatic differentiation methods.
@@ 5 Computation of the curvature forms
We report the lemmas needed to compute the tensor $N_{a,b,c,d}$ in Appendix B.
@@ 6 Experiments
[…]
@@
[…]
@@
[…]
@@ 7 Conclusion
Appendix A An auxiliary lemma
@@
Appendix B Further computations of the curvature forms

### v2

@@ 1 Introduction
In machine learning, the idea of data manifold is based on the assumption that the data space, containing the data points on which we perform classification tasks, has a natural Riemannian manifold structure. It is a quite old concept (see [1], [2] and refs therein) and it is linked to the key question of dimensionality reduction [3], which is the key to efficient data processing for classification tasks, and more. Several geometrical tools, such as geodesics, connections, Ricci curvature, become readily available for machine learning problems under the manifold hypothesis, and are especially effective when employed in examining adversarial attacks [4, 5, 6, 7, 8] or knowledge transfer questions, see [9, 10] and refs. therein.
More specifically, the rapidly evolving discipline of information geometry [11, 12, 13, 14] is now offering methods for discussing the above questions, once we cast them appropriately by relating the statistical manifold, i.e. the manifold of probability measures, studied in information geometry with data manifolds (see [11, 15]). Data manifolds have also been studied from a Riemannian geometric perspective by pullbacks of the Fisher information metric on the probability simplex through simple families of functions in the field of metric learning [16, 17].
In this paper, we are interested in a naturally emerging foliation structure on the data space coming through the data information matrix (DIM), which is the analog of the Fisher information matrix and in concrete experiments can be obtained by looking at a deep learning neural network for classification tasks. As it turns out, the leaves of such foliation are related with the dataset the network was trained with [15, 18]. Further work in [18] linked such study to the possible applications to knowledge transfer.
@@ 2 Preliminaries
$\forall i,j,k,\quad$ $\partial_{x_{i}}\partial_{x_{j}}s_{k}(x)=0,\qquad{\forall x\in\mathcal{X}\setminus A}$ (H2)
where $s$ is called a score function, $A$ is a nowhere dense set and $\mathrm{softmax}$ is the function $\left(a_{i}\right)_{i}\mapsto\left(e^{a_{i}}/\sum_{k}e^{a_{k}}\right)_{i}$. We detail why these assumptions are important in the following sections.
Examples of neural networks that satisfies these assumptions are feed-forward multi-layer perceptrons, or convolutional neural networks, with ReLU activation functions or variants such as leaky ReLU, or any piecewise linear activation functions. These kind of architectures are popular within the machine learning community.
@@
In the data information matrix, the derivative is taken with respect to the input $x$ and not with respect to the parameter $\theta$ as in the Fisher matrix. Hence, $D$ is a $d\times d$ matrix - $d$ the dimension of the dataspace - which makes it possible the study of the relationship between input-output changes in the neural network. [^1] Note that $D$ can be retrieved as the pullback of the Fisher metric on the $(C-1)$-simplex by the neural network $N$.
@@ 3 Cartan moving frames
In this section we first review some basic notions of Cartan’s approach to differentiable manifolds via the moving frames point of view and then we see some concrete application to our setting. A more detailed introduction to the Cartan moving frames approach can be found in Appendix A.
@@ 4 The Curvature of the Data Leaves
Computing in practice the curvature of a manifold with many dimensions is often intractable though essential for many tasks. This is why we are interested in computing the curvature just for the data leaves. In this section and the following one, we derive the curvature form calculations to obtain a formula linking curvature and first-order derivative of the network (i.e. the frame $e_{k}$), without double automatic differentiation or division by $p_{i}$ which might be close to zero, especially on data points, and could cause problems in practical implementation. Expressing the curvature with respect to the first-order derivative of the network can give information on the magnitude of the curvature coefficients when $N$ is almost constant in some region of the dataspace.
Notice that, since in our frame $\left(e_{k}:=\sum_{i}\partial_{i}p_{k}\left(x\right)\partial_{i}\right)$, we have $\sum_{i}\partial_{i}\ln p_{k}(x)\partial_{i}=\frac{1}{p_{k}(x)}\sum_{i}{\partial_{i}p_{k}(x)\partial_{i}}=\frac{1}{p_{k}(x)}e_{k}$, we may forget the logarithm in the computations as it contributes only by a scalar factor.
@@
The proof is a straightforward combination of Equation 10 and Equation 11. We simply need to replace $\partial_{i}p_{a}$ twice in the expression 11 by 10. ∎
@@
Then, we use the following proposition to remove the second order derivative in $M$. An alternative formula for this expression is given in B.1. This facilitates the computations with automatic differentiation methods.
@@ 5 Computation of the curvature forms
We report the lemmas needed to compute the tensor $N_{a,b,c,d}$ in Appendix C.
Remark 5.4.
The practical computation of the curvature forms involves some matrix multiplication, inversion and the computation of the first order derivatives of the network which can be done with automatic differentiation [25]. The most time consuming operation is the differentiation of the network as it scales with the number of layers (thus the complexity of the task to learn), with the input dimension and with the number of classes. Improving the efficiency of this computation is left as future work.
@@ 6 Experiments
Note that in this section, we focus on the DIM and the frame $\left(e_{a}\right)_{a=1}^{C-1}$. We do not inspect the connection or curvature forms and leave them for future work.
@@
No. 0 1 2 3 4 5 6 7 8 9
Class airplane automobile bird cat deer dog frog horse ship truck
@@
No. Layers (sequential)
(0): Conv2d(1, 32, kernel_size=(3, 3), stride=(1, 1))
@@
No. Layers (sequential)
(0): Conv2d(3, 64, kernel_size=(3, 3), stride=(1, 1), padding=(1, 1))
@@
More experiments can be found in Appendix D with full details to ensure reproducibility.
@@ 7 Conclusion
Data availability
Data sets used in this article can be found online at:
We used Python and PyTorch to access these data.
Appendix A Cartan moving frames
In this appendix we briefly review the notion of moving frames for differentiable manifolds, highlighting the key geometric properties, with some pictures to help the intuition. For more details see [23] and [26].
The theory of Cartan moving frames (repère mobile) stems from the Frenet– Serret frame theory, which lead to the celebrated Frenet-Serret formulas, developed independently by Frenet and Serret, in the XIX century, to describe the motion of a particle on a curve in $\mathbb{R}^{3}$. In fact, given smooth curve $\gamma$ in $\mathbb{R}^{3}$, a Frenet-Serret frame is an orthonormal basis of $\mathbb{R}^{3}$ attached at each point of $\gamma$ and consisting of the (normalized) tangent, normal and conormal vectors (see Fig. 3), obtained via subsequent derivations of the curve $\gamma$.
Figure 3: Frenet-Serret Frames
Notice that, once we give the curve $\gamma$, the definition of its Frenet-Serret frame is independent from the coordinates of $\mathbb{R}^{3}$; it is in fact an “intrinsic” reference frame for $\mathbb{R}^{3}$, associated to each point of the curve.
The notion of moving frames is a far reaching generalization of Frenet-Serret frames, where we replace the smooth curve $\gamma$, with a smooth manifold $M$.
Definition A.1.
We define a moving frame on a differentiable manifold $M$ a collection of vector fields $e_{1},e_{2},\dots,e_{n}$, which are a basis of $T_{p}M$, the tangent space at each point $p\in M$. (see Fig. 4). We also ask such collection to be varying smoothly with the point $p$ in $M$. The assignment of a moving frame at each point $p\in M$ is called a soldering form $\theta:T_{p}M\longrightarrow\mathbb{R}^{n}$.
Hence, once a soldering form is given, we have an intrinsic point of view on the geometry of a manifold and we can, in some sense, effectively replace the notion of local coordinates. Such point of view was successfully exploited to describe, in a local coordinate free manner, some physical theories, most notably general relativity, see [27].
In Riemannian geometry, we take the moving frames to be orthonormal frames, that is, at each point $p\in M$ the frame is an orthonormal basis of $T_{p}M$, with respect to the metric on the manifold.
Figure 4: Soldering form on a sphere.
In our application, that is the geometry of the data manifold $\mathcal{L}$, such moving frame comes naturally from the neural network and it is given by the gradients of the probability output, so that we immediately have a soldering form on $\mathcal{L}$ as follows:
$\theta:T_{p}\mathcal{L}\longrightarrow\mathbb{R}^{C-1},\qquad e_{i}\mapsto\nabla p_{i}$ (19)
Moreover we have a natural Riemannian metric on the data leaf via the DIM (data information matrix). The soldering form defines uniquely (Prop. 1.1 pg 378 in [26]) a matrix of one forms $\omega=(\omega^{i}_{j}$), called (evocatively) spin connection. So we can actually view the formula in Def. 3.4 in two equivalent ways:
  1. 1.
    Given our soldering form $\theta$ on the Riemannian manifold $\mathcal{L}$ as in (19), we have a spin connection $\omega$ which gives the Levi Civita connection by the formula (see (3) in the main text):
    $\nabla_{X}e_{j}=\sum_{i}\omega_{j}^{i}(X)e_{i}$
  2. 2.
    Given the metric via the DIM on the data leaf, we can compute its (unique) Levi Civita connection and this will give us $\omega$ the spin connection, via the above formula.
We choose to take the second point of view in our paper, to avoid introducing too much terminology and results in moving frames theory, since we prefer to rely on the correspondence between Levi-Civita connections and metrics, possibly more familiar to most readers.
These equivalent ways to view the Levi Civita connection is an illustration of the power of the moving frames method as a coordinate independent way to approach Riemannian geometry. We can indeed compute the curvature form as:
as we reported in (15) in the main text. Such curvature form is linked with the usual Riemannian curvature tensor via the equation:
$R(X,Y)e_{j}=\sum\Omega_{j}^{i}(X,Y)e_{i}$
(see [26] pg 379).
For more details on the method of moving frames and its link with the equivalent theory of Riemannian geometry making use of local coordinates, we invite the reader to consult [26] App. B.
Appendix B An auxiliary lemma
@@
Appendix C Further computations of the curvature forms
@@
Appendix D More experiments
In this section, we report some more experiments to ensure reproducibility. The code used to compute the DIM is available at https://github.com/eliot-tron/curvcomputenn/. The seed used for the following experiments is 42. On Table 4 and Table 5, we report the extreme values of the matrix $\hat{D}(x)=\left(g^{D}_{x}\left(e_{a},e_{b}\right)\right)_{a,b=1,\ldots,C}$ evaluated on six points of the test set of MNIST (see Figure 5) and CIFAR10 (see Figure 6) respectively.
Table 4: Extreme values of the matrix $\hat{D}(x)=\left(g^{D}_{x}\left(e_{a},e_{b}\right)\right)_{a,b=1,\ldots,C}$ evaluated on MNIST.
Highest $g(e_{a},e_{b})$ Second highest $g(e_{a},e_{b})$ Lowest $g(e_{a},e_{b})$
Point indices value indices value indices value
$x$ $(a,b)$ $\hat{D}(x)_{a,b}$ $(a,b)$ $\hat{D}(x)_{a,b}$ $(a,b)$ $\hat{D}(x)_{a,b}$
n°0 (4, 4) 4.4510e-11 (9, 9) 4.1848e-11 (4, 9) and (9, 4) -4.3158e-11
n°1 (7, 7) 3.9485e-05 (3, 3) 3.3201e-05 (7, 3) and (3, 7) -3.6186e-05
n°2 (9, 9) 1.7384e-04 (8, 8) 1.4993e-04 (9, 8) and (8, 9) -1.6140e-04
n°3 (3, 3) 3.9730e-08 (5, 5) 3.3480e-08 (3, 5) and (5, 3) -3.6456e-08
n°4 (8, 8) 1.8038e-08 (5, 5) 1.4309e-08 (5, 8) and (8, 5) -1.6045e-08
n°5 (2, 2) 4.1834e-09 (8, 8) 3.5215e-09 (2, 8) and (8, 2) -3.8379e-09
[…]
Figure 5: Couples of input point $x$ (above) and the corresponding matrix $\hat{D}(x)=\left(g^{D}_{x}\left(e_{a},e_{b}\right)\right)_{a,b=1,\ldots,C}$ (below) on MNIST used for Table 4 numbered from left to right and from top to bottom.
Table 5: Extreme values of the matrix $\hat{D}(x)=\left(g^{D}_{x}\left(e_{a},e_{b}\right)\right)_{a,b=1,\ldots,C}$ evaluated on CIFAR10.
Highest $g(e_{a},e_{b})$ Second highest $g(e_{a},e_{b})$ Lowest $g(e_{a},e_{b})$
Point indices value indices value indices value
$x$ $(a,b)$ $\hat{D}(x)_{a,b}$ $(a,b)$ $\hat{D}(x)_{a,b}$ $(a,b)$ $\hat{D}(x)_{a,b}$
n°0 (6, 6) 8.9736e-01 (2, 2) 8.9682e-01 (6, 2) and (2, 6) -8.9709e-01
n°1 (8, 8) 1.7268e-25 (0, 0) 6.2142e-26 (8, 0) and (0, 8) -1.0334e-25
n°2 (9, 9) 4.4503e-15 (1, 1) 4.4414e-15 (1, 9) and (9, 1) -4.4458e-15
n°3 (7, 7) 1.1507e-23 (1, 1) 1.0181e-23 (7, 1) and (1, 7) -1.0822e-23
n°4 (3, 3) 8.8027e-02 (4, 4) 6.0344e-02 (4, 3) and (3, 4) -7.2665e-02
n°5 (0, 0) 4.0292e-17 (1, 1) 3.1293e-17 (1, 0) and (0, 1) -3.5203e-17
[…]
Figure 6: Couples of input point $x$ (above) and the corresponding matrix $\hat{D}(x)=\left(g^{D}_{x}\left(e_{a},e_{b}\right)\right)_{a,b=1,\ldots,C}$ (below) on CIFAR10 used for Table 5 numbered from left to right and from top to bottom.